OpenStudy (anonymous):

WILL MEDAL AND FAN: Julie ran 32.4 miles in May and 35.8 miles in June. Greg ran the same total distance as Julie, but only ran 29.6 miles in May. How far did Greg run in June? A. 2.8 miles B. 6.2 miles C. 38.6 miles D. 33.0 miles

OpenStudy (mathmath333):

julia total run distance=may+june=32.4+35.8=68.2 greg total run distance=may+june=29.6+x Greg ran the same total distance as Julie therefore 29.6+x=68.2

OpenStudy (anonymous):

Oh okay then, x=

OpenStudy (igreen):

Julie's total distance is 32.4 + 35.8 = 68.2 Since Greg ran the same total distance, and he only ran 29.6 miles in may, he ran 38.6 miles in June. Because 68.2 - 29.6 = 38.6 miles.

OpenStudy (anonymous):

38.6

OpenStudy (igreen):

Yup.

OpenStudy (anonymous):

okay so who do I medla now? XD

OpenStudy (anonymous):

thank you my wonderful mathletes